OpenStudy (anonymous):

Find the derivative of f(x) = x^2 sin(x)

OpenStudy (anonymous):

I did \[(x^2)(\cos(x))+(\sin(x))(2x)\] and since cos(x) + sin(x) I came up with \[x^2+2x+1\] But that was wrong.

myininaya (myininaya):

So no you can't do that That is not an identity And even if it was you do not have cos(x)+sin(x) you have acos(x)+bsin(x)

myininaya (myininaya):

\[x^2\cos(x)+2x \sin(x)=x(xcos(x)+2 \sin(x))\]

OpenStudy (anonymous):

Thank you!